For the first-order reaction, 2 N2O(g) ? 2 N2(g) + O2(g), what is the concentration of N2O after 3

half-lives if 0.15 mol of N2O is initially placed into in a 1.00-L reaction vessel?
Answer #1

after 1st half life , concentration of N2O = 0.15/2 = 0.075M

after 2nd half life , concentration of N2O = 0.075/2 = 0.0375 M

after 3rd half life , concentration of N2O = 0.0375/2 = 0.01875 M
Answer #2

For first order rxn

N20 final=N20 initial e-kt

t=3ln2/k

N20 final=N20 initial/8=0.15/8= 0.01875 moles left